Lee is a cadet at the Honolulu Police Academy. He was asked to make a P-Chart for reported (minor) property crimes. Lee chose a small neighborhood with 481 families. Each family is viewed as a binomial trial. Success means that the family was the victim of at least one minor property crime in the past 3 months. Police reports gave the following data for the past 12 quarters (4 years). Assume the 481 families lived in the neighborhood all 4 years.

Quarter 1 2 3 4 5 6
r = no. of successes 11 14 18 23 19 15
= r / 481 0.02 0.03 0.04 0.05 0.04 0.03
Quarter 7 8 9 10 11 12
r = no. of successes 12 16 13 22 24 19
= r / 481 0.02 0.03 0.03 0.05 0.05 0.04

Make a P-Chart. (Use 4 decimal places.)

Center line =  Incorrect: Your answer is incorrect.
–2.0 SL =
2.0 SL =
–3.0 SL =
3.0 SL =

List any out-of-control signals by type (I, II, or III). (Select all that apply.)

Out-of-control signal I occurs on Quarter 11.There are no out-of-control signals.Out-of-control signal III occurs on Quarter 1.Out-of-control signal III occurs on Quarter 7.
Answer #1
number of samples m= 12.00
sample size n= 481.00
total number of units =mn= 5772.00
number of defects d = 206.00
hence non conforming fraction p̅=d/(mn)= 0.0357

standard error =√(p*(1-p)/n) =0.0085

control line (CL) = p̅= 0.0357

-2.0SL = 0.0357-2*0.0085 =0.0188

2SL =0.0357+2*0.0085 =0.0526

-3.0SL = 0.0357-3*0.0085 =0.0103

3.0SL =0.0357+3*0.0085 =0.0611

.There are no out-of-control signals